You purchased 18 August 13 futures contracts on soybeans at a price quote of 1,059′6. Each contract is for 5,000 bushels with the price quoted in cents and 1/8 ths of a cent per bushel. Assume the contract price is 1,069′4 when you close out your contract six weeks from now. What will be your total profit or loss on this investment?
Answer #1

price fluctuation is (1069'4-1059'6) cents= 10'2 cents= 10.25 cents= $0.1025

So, profit= 0.1025*5000*13=$6662.5 on this investment.